An Atom has 91 protons and 140 neutrons: A) Fine the mass number and the atomic number B) Find the number of electrons C) Write the hyphen notation and nuclear symbol D) Find the total number of subatomic particle.
1
Expert's answer
2016-11-26T08:28:10-0500
A) Atomic number = number of protons = 91

Mass number = number of protons + number of neutrons = 91 + 140 = 231

B) Number of electrons = Atomic number = number of protons = 91

C) Hyphen notation – Protactinium-91

Nuclear symbol - Pa

D) Total number of subatomic particles = electrons + protons + neutrons = 91 + 91 + 140 = 322